i really need help with this problem can you please help me
5. Function f(x)=ax˛+bx+c, and a>0, what happens to the graph of f(x) when:
A. a increases? B. a decreases?
C. b increases? D. b decreases?
E. c increases? F. c decreases?

Answer by stanbon(48535)

5. Function f(x)=ax˛+bx+c, and a>0, what happens to the graph of f(x) when:
A. a increases?
Compare y=x^2 to y=10x^2; notice how increasing a makes the parabola slimmer and
steeper.
-----------------
B. a decreases?
Just the opposite of the answer to A
----------------
C. b increases?
Compare y= x^2+x to y=x^2+10x; notice how the vertex moves down and to the left.
D. b decreases?
Just the opposite of the answer to C
---------------------
E. c increases?
Compare y=x^2+1 to x^2+10; notice how all the points move up
F. c decreases?
Notice how all the points move down.
-----------------
Cheers,
Stan H.